OpenStudy (anonymous):

find the laplace transform of the function f(t) = 3e^-4t + e^2t + 4t^2 - 5 find the laplace transform of the function f(t) = 3e^-4t + e^2t + 4t^2 - 5 @Mathematics

OpenStudy (anonymous):

F(s)=(3/(s+4))+(1/(s-2))+8/(s^3)-5/s but not 100% sure,there are formulas for the laplace transform of functions..

OpenStudy (anonymous):

yeah that is what i got but i am unsure about simplification? is that an adequate final answer?

OpenStudy (anonymous):

sure it is if u take a common denominator you'll make just more complicated

OpenStudy (anonymous):

yeah fair enough cool cheers dude

OpenStudy (anonymous):

what happens when you have a product? do you do the laplace seperately then multiply the two together?

OpenStudy (anonymous):

no if not to find in the tables u have to use the formula F(s)=\[\int\limits_{0}^{\infty}f(t)e ^{-st}dt\]

OpenStudy (anonymous):

g(t) = te^(2t)

OpenStudy (anonymous):

tf(t) gives in the s domain -F'(s),where F(s) is(1/(s-2)) ,derive and you;ll get the same answer the website gives..

OpenStudy (anonymous):

i know next to nothing about this, but here is a very basic and excellent introduction courtesy of jamesj http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-19-introduction-to-the-laplace-transform/